OpenStudy (anonymous):

Find the centre of mass of three particle at the vertices of an equilateral triangle. The masses of the particles are 100g , 150g and 200 g.Each side of the equlateral triangle is 0.5m long?

OpenStudy (ghazi):

(m1x1+m2x2+m3x3)/ (x1+x2+x3)

OpenStudy (ghazi):

or else you can use double integral to find center of mass easily...

OpenStudy (anonymous):

hw??

OpenStudy (anonymous):

(m1x1+m2x2+m3x3)/ (m1 + m2 + m3)

OpenStudy (anonymous):

good question

OpenStudy (ghazi):

 now just define a function that passes through the line of negative slope ....and decide the limits...you'll have center of mass

OpenStudy (anonymous):

hw to find using (m1x1+m2x2+m3x3)/ (m1 + m2 + m3)

OpenStudy (anonymous):

x1 = 0.5????

OpenStudy (ghazi):

\[\int\limits_{0}^{0.5}\int\limits_{0}^{-0.25x} dxdy\]

OpenStudy (anonymous):

not intergrating...

OpenStudy (anonymous):

i have nt studied...much abt this

OpenStudy (anonymous):

hw to find using (m1x1+m2x2+m3x3)/ (m1 + m2 + m3)

OpenStudy (ghazi):

or else you can use X= (100*0.5+150*0.5+200*0.5)/(100+150+200)...similarly for Y

OpenStudy (anonymous):

y = 0.5??

OpenStudy (ghazi):

one more thing that you can do is ....consider the equilateral triangle that i am showing you now...and find X and Y